OpenStudy (anonymous):

sin B= 0.8 What is the length of AB?( right triangle ABC has a base, AC, of 12 with A on the left point and C on the right point of the base. AB is the hypotenuse and BC is the side.)

OpenStudy (alexwee123):

didn't i help you out with a problem like this like 5 minutes ago?

OpenStudy (alexwee123):

since this is sin remember that sin=opposite over hypotenuse

OpenStudy (alexwee123):

sine*

OpenStudy (anonymous):

i thought it looked familiar.

OpenStudy (anonymous):

lol yeah but i got stuck with a different one because it has a different measurement. I answered one on my own with your help from that problem i posted.

OpenStudy (anonymous):

i think its answer is 15

OpenStudy (arnavguddu):

sin B = height/hyp = AC/AB but u know base, so find AC first. Then, use pythagoras theorem, AB2=AC2+BC2 from this, BC2=AB2-AC2. take square root of bc2, and u will find bc, try it and reply????

OpenStudy (anonymous):

I got 14.976(15) would that be right? I turned .8 into a fraction 1.25/1.56=12/x= 18.72/1.25=15